If dim V < infinity and T is a linear transformation such that for all v belongs to V , there exist a positive integer k such that(T)^k(v) =0, then prove that T is nilpotent.
1
Expert's answer
2021-07-20T17:29:28-0400

Let "\text{dim} \ V=n" and let "\ T:\ V\rightarrow V" be a linear transformation such that for all "v\in V," there exists integer "k>0" such that "T^k v=0" .

Suppose "\beta =\{\beta_1,\ \beta_2,\ …\ , \ \beta_n\}" is a basis for "V" .

For each vector "\beta _i" there exists "k_i" such that "T^{k_i}\beta_i=0" .

Let "k=\max \{ k_1,\ k_2,\ … , \ k_n\}" . In this case we have that "T^k\beta _i=T^{k-k_i} \left(
T^{k_i}\beta_i
\right)=T^{k-k_i}(0)=0" .

Since every vector "v\in V" is a linear combination of "\beta_i" , we can consider "v=\sum\limits_{i=1}^nc_i\beta_i" for some "c_i" .

"T^kv=T^k \left(\sum\limits_{i=1}^nc_i\beta_i\right)= \sum\limits_{i=1}^nc_i T^k(\beta_i)=\sum \limits_{i=1}^nc_i\cdot 0=0" .

So, "T^k=0" and it proves that "T" is nilpotent.